# Hybrid Deepfake Image Detection: A Comprehensive Dataset-Driven Approach Integrating Convolutional and Attention Mechanisms with Frequency Domain Features

### v1

@@ Abstract
Effective deepfake detection tools are becoming increasingly essential over the last few years due to the growing usage of deepfakes in unethical practices. There exists a diverse range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism. The 2025 Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images, which are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection. To this end, we proposed an ensemble-based approach that employs three different neural network architectures: a ResNet-34-based architecture, a data-efficient image transformer (DeiT), and an XceptionNet with Wavelet Transform to capture both local and global features of deepfakes. We visualize the specific regions that these models focus for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots. Individually, the ResNet-34 architecture has achieved 88.9% accuracy, whereas the Xception network and the DeiT architecture have achieved 87.76% and 89.32% accuracy, respectively. With these networks, our weighted ensemble model achieves an excellent accuracy of 93.23% on the validation dataset of the SP Cup 2025 competition. Finally, the confusion matrix and an Area Under the ROC curve of 97.44% further confirm the stability of our proposed method.
@@
Deepfake , ResNet-34 , Transformer , DeiT , Wavelet , XceptionNet
journal: Elsevier Image and Vision Computing
@@ 1 Introduction
Deepfakes are media that are edited or created by generative Artificial Intelligence (AI), which do not exist or have not occurred in reality. Although deepfakes have existed for a long time, it took experts, and often entire studios, to generate convincing fake media. Due to the surge of machine learning and generative AI tools, the generation of deepfakes has never been easier and more accessible. Such ease of access has raised serious societal concerns since these are being used to manipulate public opinion, commit fraud, create non-consensual pornography, and attack individuals through defamation or identity threats [1]. Deepfakes are also being used to gain malicious political advantage worldwide [2]. Although deepfake generation tools are widely accessible, detection tools are not nearly as effective and accurate. More specifically, the existing detection methods only work well when the test image resembles the features of the images of the training dataset of the detector. That is, the major issues in the existing detection approaches are the lack of diversity in training datasets, and the inability to be up to date with new fake image generation techniques [3].
To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm. Such a diverse dataset poses unique challenges to formulating accurate image detection methods. In our proposed solution for deepfake image detection, we have employed a weighted ensemble approach, which utilized three different neural network architectures with different strengths: i) a ResNet-34-based architecture with squeeze and excitation block, ii) a data-efficient image transformer (DeiT) model, and iii) an Xception network. More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result. We note that the ResNet-34-based architecture, due to its residual connections, allows learning complex hierarchical features with deeper networks by addressing the problem of vanishing gradients [4]. Moreover, the squeeze and excitation (SE) block incorporated in the ResNet-34-based architecture provides channel-wise attention with very little computational expense [5]. Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image. More specifically, the ResNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [6]. Finally, the Xception model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images. We have utilized the ’Haar’ wavelets due to their notable ability to detect abrupt transitions and edges in the data [7]. A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone.
@@
Related Works. The advancement in deepfake research has led to the creation of nearly perfect image manipulations that are undetectable to the human eye. The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [8] showed ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [9] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset. Wolter et al. [10] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [11] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [12]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [13]. However, Abdullah et al. [14] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [15, 16]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [17]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [18]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [18]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [19, 20].
@@ 2.1 Description of the Dataset
As mentioned before, the DFWild-Cup competition has provided a diverse dataset. The dataset contains images from eight publicly available standard datasets designed for the DeepfakeBench evaluation [25]: Celeb-DF-v1 [26], Celeb-DF-v2 [26], FaceForensics++ [27], DeepfakeDetection [28], FaceShifter [29], UADFV [30], Deepfake Detection Challenge Preview [31], and Deepfake Detection Challenge [32]. To ensure that the detector performance is not biased by specific pre-processing techniques, all datasets underwent uniform pre-processing. Additionally, any information that would identify the origin of the dataset has been eliminated, and the file names have been anonymized. For training, there are a total of 42,690 real images and 219,470 fake images. The validation dataset consists of 1,548 real images and 1,524 fake images. It is evident that there is a significant class imbalance present in the training data as the ratio of real and fake images is approximately 5:1.
Figure 1: Training procedure including class imbalance handling
@@ 2.2.1 Class Balancing for Model Training
Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, was trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
@@
We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [33]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, following the optimization process:
@@
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images. Figure 1 illustrates the sequential learning procedure.
2.2.2 Data Augmentation
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips, random rotations, color jittering, and random resized cropping. The random horizontal flip method flips the training images along their vertical axis with a probability of 50%. The Random rotation method augments images by randomly rotating them within $0^{o}-10^{o}$. The color jitter method randomly alters the brightness, contrast, saturation, and hue of images within a specified range. To this end, a range of 0.2 for brightness, contrast, saturation, and hue provided the best results in our experiments. Finally, the random resized crop was used as a data augmentation technique that applies both random cropping and resizing to create varied versions of training images. In our case, it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the ResNet-34 and the DeiT architectures. For the XceptionNet, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain.
2.2.3 Data Resizing and Normalization
All three models were pre-trained using the ImageNet dataset. Since the requirements of input image dimensions are different on the ResNet-34 architecture and the Xception network, we have resized the input images accordingly to the specific dimensions for respective architectures. Additionally, the training images were normalized using the mean and standard deviation of the ImageNet dataset [33].
2.3 Wavelet Feature Extraction for XceptionNet
@@
Figure 2: Row and Column-wise filtering and downsampling to generate approximation and detail images through Haar 2D-DWT
@@
To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the XceptionNet model. In Figure 3, we showed two such feature images with the approximate and detailed coefficients of two sample images from the dataset.
@@ 2.4 Details of the Model Architectures
2.4.1 ResNet-34 with SE Block
We implemented a ResNet-34 model with SE block for the detection of deepfake images. Note that, we investigated with different Convolutional Neural Network (CNN)-based architectures – shallow CNNs exhibited high bias, whereas deep CNNs caused vanishing gradients. The residual mapping of the ResNet-34 architecture addresses the issue of vanishing gradients, while the capacity of the ResNet-34 architecture is sufficient for capturing the intricate features for deepfake classification.
Figure 4: Building block of residual learning
Note that residual learning works by reformulating the target function $\text{H}(x)$ as $\text{F}(x)=\text{H}(x)-x$, where $x$ is the input [4]. Instead of directly learning $\text{H}(x)$, the network learns the residual $\text{F}(x)$. The final output is computed as $\text{F}(x)+x$, combining the learned residual with the input via a skip connection. This mechanism allows each layer to focus on learning the residual adjustments needed to improve upon the input. The building block of residual mapping is shown in Figure 4. Our ResNet-34-based architecture is implemented with two simple principles following the traditional structure of this model: i) When maintaining the same output feature map size, each layer will consistently have the same number of filters. ii) Additionally, when the feature map size is halved, the number of filters is effectively doubled. Based on these principles, shortcut connections are implemented on a generic CNN architecture to obtain the desired residual network.
Figure 5: ResNet-34 architecture with Squeeze and Excitation
Squeeze and Excitation (SE) Block. The SE block is introduced as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [5]. The SE block is placed after the final sequential layer (layer 4) of the ResNet-34 model to enhance feature representation by modeling channel-wise dependencies. It is responsible for two main operations:
  • 1.
@@ 2.4.1 ResNet-34 with SE Block
  • 2.
    Excitation: The descriptors generated from the Squeeze block are passed through a lightweight two-layer fully connected network with a bottleneck structure to model non-linear channel independencies. The squeezed feature vector is passed through a gating mechanism with two fully connected layers and a non-linearity (typically ReLU and Sigmoid) to learn channel-wise dependencies:
    $s=\sigma(W_{2}\delta(W_{1}z)),$ (7)
    where $W_{1}\in\mathbb{R}^{\frac{C}{r}\times C}$ and $W_{2}\in\mathbb{R}^{C\times\frac{C}{r}}$ are the weights of the two fully connected layers, $\delta(\cdot)$ represents the ReLU activation function, and $\sigma(\cdot)$ is the Sigmoid function that outputs a channel-wise scaling factor $s\in\mathbb{R}^{C}$. The final output is obtained by rescaling the input feature maps $X$ using the channel-wise attention weights:
    $\tilde{X}_{c}=s_{c}X_{c},\quad c=1,2,...,C$ (8)
    This step selectively enhances informative channels while suppressing less important ones. We denoted this operation as a function $F_{ex}(.)$, as depicted in Fig. 5.
The details of the ResNet-34 architecture including the output shapes of each layer and the number of parameters are shown in Table 1.
Table 1: Details of the ResNet-34 Architecture with SE Block
@@
Conv2d, BatchNorm2d, ReLU [64, 112, 112] 9,536
MaxPool2d [64, 56, 56] 0
BasicBlock-1 $\times$ 3 [64, 56, 56] 221,952
   —Conv2d [64, 56, 56] 36,864
   —BatchNorm2d + ReLU [64, 56, 56] 128
   —Conv2d [64, 56, 56] 36,864
   —BatchNorm2d + ReLU [64, 56, 56] 128
BasicBlock-2 $\times$ 4 [128, 28, 28] 1,116,416
BasicBlock-3 $\times$ 6 [256, 14, 14] 6,822,400
BasicBlock-4 $\times$ 3 [512, 7, 7] 13,114,368
SEBlock (Global Avg Pool + FC) [512, 7, 7] 33,312
AdaptiveAvgPool2d [512, 1, 1] 0
Linear [2] 1,026
Total Parameters 21,319,010
Trainable Parameters 8,204,642
Non-Trainable Parameters 13,114,368
2.4.2 Data Efficient Image Transformer (DeiT) Model
Unlike conventional Recurrent Neural Networks (RNNs) that process sequential data step by step, transformers process all inputs simultaneously, relying on a self-attention mechanism. This simultaneous processing allows transformers to learn long-range dependencies and global contexts of data [6]. It has been shown that the data-efficient image transformer, which is based on the backbone of the vision transformer, works well for cases where the training data is limited [36]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [37]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$ and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}.$

### v2

# CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features

@@ Abstract
Effective deepfake detection tools are becoming increasingly essential to the growing usage of deepfakes in unethical practices. There exists a wide range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism. The 2025 IEEE Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images containing significant class imbalance. The images in the dataset are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection. To this end, we proposed a disjoint set-based multistage training method to address the class imbalance and devised an ensemble-based architecture CAE-Net. Our architecture consists of a convolution- and attention-based ensemble network, and employs three different neural network architectures: EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet transform to capture both local and global features of deepfakes. We visualize the specific regions that these models focus on for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots. Individually, the EfficientNet B0 architecture has achieved 90.79% accuracy, whereas the ConvNeXt and the DeiT architecture have achieved 89.49% and 89.32% accuracy, respectively. With these networks, our weighted ensemble model achieves an excellent accuracy of 94.63% on the validation dataset of the SP Cup 2025 competition. The equal error rate of 4.72% and the Area Under the ROC curve of 97.37% further confirm the stability of our proposed method. Finally, the robustness of our proposed model against adversarial perturbation attacks is tested as well, showing the inherent defensive properties of the ensemble approach.
@@
Deepfake detection , ConvNext , Transformer , DeiT , Wavelet , EfficientNet
@@ 1 Introduction
Deepfakes are media that are edited or created by generative Artificial Intelligence (AI), which do not exist or have not occurred in reality. Although deepfakes have existed for a long time, it took experts, and often entire studios, to generate convincing fake media. Due to the surge of machine learning and generative AI tools, the generation of deepfakes has never been easier and more accessible. Such ease of access has raised serious societal concerns since these are being used to manipulate public opinion, commit fraud, create non-consensual pornography, and attack individuals through defamation or identity threats [1]. Deepfakes are also being used to gain malicious political advantage worldwide [2]. Although deepfake generation tools are widely accessible, detection tools are not nearly as effective and accurate. Thus, effective detection of deepfakes is a task of great importance and has been extensively studied for the past decade.
@@
The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [4] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset. Wolter et al. [5] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [6] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [7]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [8]. However, Abdullah et al. [9] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [10, 11]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [12]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [13]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [13]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [14, 15].
@@
Research gap and our contribution. The existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector. That is, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features and as such, ii) the inability to be up to date with new fake image generation techniques [20]. To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm. Such a diverse dataset poses unique challenges to formulating accurate image detection methods. In our proposed solution for generalized deepfake image detection, we have employed a weighted ensemble approach, CAE-Net, which utilizes three different neural network architectures with different strengths: i) an EfficientNet architecture, ii) a Data Efficient Image transformer model, and iii) a ConvNeXt model. More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result. We note that the EfficientNet-based architecture, due to its depthwise separable convolution and squeeze and excitation block, allows learning fine spatial artifacts and enhances channel-wise attention [21]. Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image. More specifically, the EfficientNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [22]. Finally, the ConvNeXt model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images. We have utilized the “Haar” wavelets due to their notable ability to detect abrupt transitions and edges in the data [23]. A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone. Finally, we test our models against adversarial perturbation attacks, which shows the trade-offs between weighted ensemble and majority voting ensemble methods.
The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, the description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt. Section 3 contains the accuracy results of multiple standard models on the dataset, along with the implementation details and other evaluation metrics of our proposed approach. This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations. Finally, Section 4 draws concluding remarks on the contributions presented in the paper.
@@ 2.1 Description of the Dataset
As mentioned before, the DFWild-Cup competition has provided a diverse dataset. The dataset contains images from eight publicly available standard datasets designed for the DeepfakeBench evaluation [24]: Celeb-DF-v1 [25], Celeb-DF-v2 [25], FaceForensics++ [26], DeepfakeDetection [27], FaceShifter [28], UADFV [29], Deepfake Detection Challenge Preview [30], and Deepfake Detection Challenge [31]. To ensure that the detector performance is not biased by specific pre-processing techniques, all datasets underwent uniform pre-processing. Additionally, any information that would identify the origin of the dataset has been eliminated, and the file names have been anonymized. For training, there are a total of 42,690 real images and 219,470 fake images. The validation dataset consists of 1,548 real images and 1,524 fake images. It is evident that there is a significant class imbalance present in the training data, as the ratio of fake and real images is approximately 5:1.
@@ 2.2.1 Class Balancing for Model Training
One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset. Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, are trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
@@
We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [32]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, through the optimization:
@@
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images.
2.2.2 Data Augmentation and Normalization
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips(with 50% probability), random rotations(within $0^{o}-10^{o}$), color jittering (20% variation in brightness, contrast, saturation and hue), and random resized cropping where it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the DeiT architectures. For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain. Since all three models were pre-trained using the ImageNet dataset, the training images were normalized using the mean and standard deviation of the ImageNet dataset [32]. The input images were resized accordingly to the specific dimensions for the respective architectures.
2.3 Wavelet Feature Extraction for ConvNeXt
@@
Figure 1: Row- and column-wise filtering and downsampling to generate approximation and detail images through 2D-Discrete Wavelet Transform (DWT) using Haar wavelets.
@@
To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the ConvNeXt model. In Figure 2, we showed two such feature images with the approximate and detailed coefficients of two sample images from the dataset.
@@ 2.4 Details of the Model Architectures
2.4.1 EfficientNet
EfficientNet-B0 is a computationally efficient model whose architecture is optimized via neural architecture search (NAS) to maximize accuracy under a FLOP constraint, resulting in a highly efficient baseline that outperforms models with similar computational cost [21]. The core block of the EfficientNet architecture is the Mobile Inverted Bottleneck Convolution (MBConv) block inspired by MobileNetV2 [35]. The MBConv block consists of a 1x1 Convolution, which expands the input channels to increase feature dimensionality. Then, a 3x3 or 5x5 depthwise convolution is applied to capture spatial features efficiently. Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [36]. The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [37]. It is responsible for two main operations:
@@ 2.4.1 EfficientNet
Excitation: The descriptors generated from the Squeeze block are passed through a lightweight two-layer fully connected network with a bottleneck structure to model non-linear channel independencies.
After that, a skip connection adds the input to the output if the input and output shapes match. The swish activation [38] function is applied after each convolution except the final projection.
$\text{Swish}(x)=x\cdot\sigma(x)=\frac{x}{1+e^{-x}}$ (7)
EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage. A brief summary of different layers, the output shapes and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
Table 1: Summary of the EfficientNetB0 Model Architecture
@@
Conv2D + BatchNorm + SiLU [32, 112, 112] 928
DepthwiseSeparableConv [16, 112, 112] 1,448
InvertedResidual-1 $\times$ 2 [24, 56, 56] 16,714
InvertedResidual-2 $\times$ 2 [40, 28, 28] 46,640
InvertedResidual-3 $\times$ 3 [80,14,14] 242.930
InvertedResidual-4 $\times$ 3 [112, 14, 14] 543,148
InvertedResidual-5 $\times$ 4 [192, 7, 7] 2,026,348
InvertedResidual-6 [320, 7, 7] 717,232
Conv2d + BatchNorm + SiLU [1280, 7, 7] 412,160
AdaptiveAvgPool2d + Flatten [1280] 0
Linear [1] 1,281
Total Parameters 4,008,829
Trainable Parameters 4,008,829
Non-Trainable Parameters 0
2.4.2 Data Efficient Image Transformer (DeiT)
Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism. This allows transformers to learn long-range dependencies and global contexts of data [22]. The data-efficient image transformer, based on the backbone of the vision transformer, has been shown to work well in cases where training data is limited [39]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [40]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$, and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}$ (8)